write an equation in slop-intercept form through (3,-5) and (4,1) and thank you so very much.
Answer by mananth(16946)

The slope of line passing through (x1,y1),(x2,y2) = (y2-y1)/(x2-x1)

(3,-5) and (4,1)
m= 1+5/4-3
m=6/1
m=6
..
Find b from y = mx+b
1=6*4+b
1=24+b
b=-23
y= 6x-23